OpenStudy (anonymous):

tell whether the lines are parallel, perpendicular, or neither. line 1 (-1,9),(-6,-6) line 2 (-7,-23), (0,-2).

OpenStudy (anonymous):

To answer this question, you have to find the slopes. slop of the line 1, m1 = (y1-y2)/(x1-x2)= (9-(-6)) / (-1-(-6)) slop of the line 2, m2 = (y1-y2)/(x1-x2)= (-23-(-2)) / (-7-0) Now you can find m1, m2. If lines are parallel, they have the same slopes. m1=m2 If lines are perpendicular, they have slopes that are negative reciprocals of each other. m1=-1/m2 If the slopes are equal or negative reciprocals they are neither parallel nor perpendicular.

OpenStudy (anonymous):

can u graph it as well??

OpenStudy (anonymous):

@sonja_lee2 Actually there's no need of graphing the two lines. because you can determine whether the lines are parallel or perpendicular using the slopes. Here, slop of the line 1, m1 = (y1-y2)/(x1-x2)= (9-(-6)) / (-1-(-6)) = 3 slop of the line 2, m2 = (y1-y2)/(x1-x2)= (-23-(-2)) / (-7-0) = 3 m1= m2 That means the two lines are parallel 

OpenStudy (anonymous):

ohh okay i see